### v1

# Low-Level and NUMA-Aware Optimization for High-Performance Quantum Simulation

Abstract
Scalable classical simulation of quantum circuits is crucial for advancing both quantum algorithm development and hardware validation. In this work, we focus on performance enhancements through meticulous low-level tuning on a single-node system, thereby not only advancing the performance of classical quantum simulations but also laying the groundwork for scalable, heterogeneous implementations that may eventually bridge the gap toward noiseless quantum computing. Although similar efforts in low-level tuning have been reported in the literature, such implementations have not been released as open-source software, thereby impeding independent evaluation and further development. We introduce an open‑source, high‑performance extension to the QuEST simulator that brings state‑of‑the‑art low‑level and NUMA optimizations to modern computers. Our approach emphasizes locality-aware computation and incorporates hardware-specific optimizations such as NUMA-aware memory allocation, thread pinning, AVX-512 vectorization, aggressive loop unrolling, and explicit memory prefetching. Experiments demonstrate significant speedups – 5.5–6.5× for single-qubit gate operations, 4.5× for two-qubit gates, 4× for Random Quantum Circuits (RQC), and 1.8× for Quantum Fourier Transform (QFT), demonstrating that rigorous performance tuning can substantially extend the practical simulation capacity of classical quantum simulators on current hardware.
Index Terms:
CPU; quantum simulation; high performance computing; NUMA; SIMD; QuEST; QFT; RQCI Introduction
Classical simulation of quantum circuits is still indispensable for validating algorithms, benchmarking hardware, and debugging control software, while error-corrected quantum computers remain beyond current reach. In the full state‑vector simulation an -qubit state is represented by complex amplitudes, so both the memory footprint (16 bytes per amplitude in double-precision) and the number of floating‑point operations grow exponentially with . Even a modest 35‑qubit circuit already requires 0.5 TiB of memory and billions of complex arithmetic operations, quickly exhausting today’s servers. Scaling beyond a few dozen qubits remains challenging because hardware and software bottlenecks now dominate performance. Memory bandwidth and capacity are primary constraints: even on multi‑socket non‑uniform memory access (NUMA) servers, accessing a ‑element state-vector saturates the memory hierarchy [1]. In classically distributed simulation of quantum computing, the network communication required for qubit gates – whether single- or two-qubit – spanning different nodes represents a huge overhead and often dominates the run time [1, 2]. At the software level, most simulators still under-use low‑level hardware features, leaving appreciable performance on the table and underscoring the need for architecture‑aware refinements.
State-of-the-art/practice. To push the simulation limits, a suite of high‑performance state‑vector simulators has emerged that exploit modern hardware architectures and parallelism [3]. QuEST [4], ProjectQ [5], Intel‑QS (qHiPSTER) [6], Qrack [7], Qulacs [8], Quantum++ [9], qsim [10] and, most recently, PennyLane Lightning [11] are representative state-of-the-art examples – compared in Table I. Among many simulators exploiting multi‑threading (OpenMP), distributed Message Passing Interface (MPI) execution, and GPU off‑loading to enlarge the range of tractable problem sizes, some state-vector simulators accelerate the core operation (applying gate matrices to the state vector) using explicit single-instruction-multiple-data (SIMD) parallelism. By operating on multiple amplitudes in one CPU instruction, SIMD vectorization can significantly speed up state updates [4]. For example, the ProjectQ simulator explicitly supports Intel AVX vector instructions [5]. More recently, Google’s qsim is built with AVX2 and fused multiply-add (FMA) intrinsics, using fused gate kernels to exploit data-level parallelism [10]. Xanadu’s PennyLane Lightning simulator similarly employs explicit intrinsic AVX-512, effectively unrolling loops to operate on 8 or 16 values at a time [11]. These approaches reduce instruction overhead by applying gate operations to multiple state-vector entries concurrently. Academic simulators have reported substantial speedups from SIMD optimizations. The HpQC simulator used AVX2/FMA instructions to maximize SIMD utilization, combined with manual bit-level optimizations [12]. This yielded about an average speedup of 2.20× (GNU compiler) over QuEST in QFT benchmarks on a multi-core CPU. The authors attribute this gain largely to vectorized gate kernels and reduced instruction count. Another effort, PAS [13], introduced a hybrid vectorization scheme alongside other tricks, achieving up to several-fold speedups versus QuEST.
| Simulator | Vectorisation | Manual Unroll | Cache Blocking | OMP | Thread Pinning | NUMA | MPI | GPU |
| ProjectQ | AVX2 (comp.) | ✗ | ✓ | ✓ | — | ✗ | ✗ | ✗ |
| qsim | AVX2 (intrin.) | ✗ | ✓ | ✓ | — | ✗ | ✗ | ✗ |
| Lightning | AVX2/512 (intrin.) | ✓ | ✗ | ✓ | likely | ✗ | ✓ | ✓ |
| HpQC/PAS | AVX2 (intrin.) | ✗ | ✗ | ✓ | — | ✗ | ✓ | ✗ |
| Qulacs | AVX2 (intrin.) | ✗ | ✗ | ✓ | — | ✗ | ✗ | ✓ |
| Qiskit Aer | AVX2 (comp.) | ✗ | ✗ | ✓ | — | ✗ | ✗ | ✓ |
| Intel‑QS | AVX512 (comp.) | ✗ | ✗ | ✓ | likely | ✗ | ✓ | ✗ |
| QuEST | AVX2 (comp.) | ✗ | ✗ | ✓ | — | ✗ | ✓ | ✓ |
| NWQ‑Sim | AVX2 (comp.) | ✗ | ✗ | ✓ | — | ✗ | ✓ | ✓ |
“Aligned Memory”, not included, is largely implicit (most tools use 64‑byte alignment for AVX‑512), while “Thread Pinning” and “NUMA‑Aware allocation” are seldom discussed—Lightning and Intel‑QS are inferred to pin threads via OMP_PROC_BIND.
comp. = compiler auto‑vectorisation; intrin. = handwritten intrinsics.
likely = affinity hinted at in code or docs but not explicitly evaluated in the paper;
✓/✗ = feature explicitly present or absent in publication, documentation or source.
To fully leverage SIMD, simulators ensure proper memory alignment of the state-vector (often 32- or 64-byte aligned for AVX2/512). Although often not explicitly discussed in publications, alignment is a standard practice to avoid penalties when loading vectors into registers. Many simulators also focus on cache-friendly memory access patterns. One common technique is cache blocking – structuring computations to maximize reuse of data in fast caches before moving to slower memory. In quantum circuit terms, this translates to fusing consecutive gate operations so that the state-vector is traversed fewer times. At a higher level, circuits can also be rearranged to minimise communication, effectively cache-blocking at the node level [14, 15]. ProjectQ and Google’s qsim use such gate fusion to improve locality. By multiplying small gate matrices into a larger effective matrix, the simulator can apply a batch of operations with one pass over the state, thus improving cache utilization and reducing memory bandwidth pressure. Beyond fusion, recent research introduced novel data layouts and filters to minimize unnecessary memory traffic. PAS describes a “memory access filtering” method, which avoids touching state-vector regions that are unaffected by a given gate. Likewise, PAS’s use of fast bitwise operations optimizes the index calculations for mapping qubit operations to memory locations, replacing expensive arithmetic or branch-heavy code with efficient bit-manipulation. The HpQC tool also emphasized a custom data structure to exploit spatial locality in the cache. In practice, these low-level memory optimizations yielded measurable speedups: e.g., HpQC reported 1.74x (GNU compiler) speed improvement on RQCs vs. baseline QuEST by cutting down redundant memory accesses.
Today’s Hardware. Modern CPUs offer dozens of cores, and virtually all state-vector simulators use multithreading to split the amplitude space across cores. OpenMP is commonly employed to parallelize the innermost loops of state updates. QuEST and others leverage multi-core parallelism by dividing the state-vector into chunks processed concurrently by threads. This yields near-linear speedup with core count [4] until memory bandwidth becomes the bottleneck [16]. However, multithreading on large shared-memory systems introduces challenges related to thread affinity and load balancing. If the threads are not carefully pinned or if the work is uneven, performance can suffer. Notably, thread pinning (assigning each thread to a specific core) is often needed to prevent Operating System (OS) scheduler overhead or migration between cores. Additionally, on multi-socket NUMA nodes, pinning helps keep threads on the socket where their data resides. While most quantum simulator papers do not explicitly detail their thread affinity policies, the need is recognized in the community [17].
On NUMA architectures, memory is physically partitioned, and accessing remote memory across the interconnect (e.g., QPI/UPI) happens transparently, i.e., without explicit communication, but incurs higher latency and lower bandwidth. It is therefore crucial for large state-vector simulations to be NUMA-aware. A straightforward strategy is to partition the state-vector across NUMA nodes and pin threads such that each NUMA domain primarily works on its portion of the state. In an ideal scenario, each thread only accesses the segment of the state-vector that resides in its local node’s memory [17]. This avoids the costly scenario where a thread continuously fetches data located in another socket’s RAM. When thread affinity and memory placement are aligned, memory bandwidth scales with the number of NUMA nodes, instead of being bottlenecked by slower cross-socket memory access [18]. Despite its importance, NUMA-aware optimization has not always been a focal point in quantum simulator research. Many open-source simulators rely on the OS’s default first-touch policy to allocate memory locally, but do not report additional NUMA-specific techniques in their publications. We found little explicit analysis of NUMA effects in simulator papers, indicating a gap that our work aims to fill. A recent IBM patent [17] describes how a naive parallelization of a CNOT gate can lead to one socket’s thread constantly accessing another socket’s memory (an imbalanced situation) and proposes reordering qubit indices or replicating data to balance the memory accesses. The fact that such patent-level solutions exist suggests that while NUMA issues are known, they have not been thoroughly addressed in open literature.
Contributions. Motivated by these observations, and building on the strong foundation of QuEST v3.7111In preparing this draft, we noted that QuEST v4 was released, introducing several software architectural enhancements – including revised data structures that partially overlap with our improvements – but it does not incorporate explicit SIMD vectorization or NUMA-aware optimizations as presented in this work. We provide further details in Section V., we introduce a suite of single‑node enhancements intended as a springboard for efficient multi‑node and GPU scaling. Our extension offers tunable modules: (i) a redesigned data layout paired with hand‑written AVX‑512 kernels; (ii) NUMA‑aware memory allocation with enforced thread pinning; and (iii) aggressive loop unrolling combined with software prefetching. Consequently, three main gaps in the literature are filled: (i) open-source availability and reproducibility of all these low-level tricks in one place – ensuring that the improvement is not a one‑off. Many prior works were either proprietary or did not release their full code (e.g., HpQC, PAS), or they focused on a single aspect (e.g., PAS on single-node CPU, Intel-QS on distributed scaling) without covering the entire optimization spectrum. (ii) The lack of fine-grained architectural tuning. For example, most simulators do their optimizations internally and automatically; there is little ability to adjust, say, toggle on/off a specific optimization for comparison. Our work provides a more tunable environment, where one can study the effect of, e.g., enabling NUMA-aware allocation vs. not, or using AVX2 vs. AVX-512. This is highly valuable for scientific computing research, as it enables understanding the contribution of each optimization. It also aids in targeting different architectures – e.g., older CPUs without AVX-512 might benefit from different tuning than the latest servers – something an explicitly tunable simulator can accommodate. Finally, (iii) the focus on NUMA-aware execution is filling a clear gap. Prior open-source tools largely assumed a uniform shared-memory model and did not optimize for multiple NUMA regions, which can leave performance on the table in multi-socket machines. By implementing strategies such as allocating state subarrays per NUMA node and pinning threads, we can significantly reduce interconnect traffic.
Paper Organization. Section II details our system architecture and optimization methodology—including our novel data structure layout (Section II-A), NUMA-aware optimization strategies (Sections II-B1 and II-B2), and various instruction-level tuning techniques (Section II-C). Section III presents our benchmark results and analysis, covering single- and two-qubit operations, QFT, and RQC performance. As demonstrated, integrating our newly designed data layout and these targeted optimizations yields performance gains comparable to those reported by HpQC and PAS, although not as pronounced as some of their most optimistic findings. Section IV discusses the scaling and portability of our framework. Finally, Section V concludes with a summary of our findings and a discussion of potential directions for future research.
II System Architecture and Methodology
II-A Data Structure Enhancements
In the baseline QuEST implementation, complex amplitudes were stored in a structure-of-arrays (SoA) format. Each state-vector was represented by a ComplexArray object containing two separate arrays: one for all real parts and one for all imaginary parts. While this representation simplifies certain operations, it leads to inefficient memory access and poor cache utilization because accessing a complex amplitude requires touching two distant memory locations. We refactored this into an array-of-structures (AoS) layout, packing each complex number’s real and imaginary components contiguously in memory. This AoS format (illustrated in Figure 1) places the real and imaginary parts of each amplitude next to each other, improving spatial locality. As a result, when iterating through the state vector, each cache line fetched contains complete complex values, reducing cache misses and better exploiting hardware prefetching. We also ensure that the allocated array of complex numbers is aligned in memory, which facilitates aligned vector load/store operations in later optimizations. This transition from an AoS to a SoA allows for improved performance in both memory usage and data access patterns, particularly in high-performance contexts like quantum computing simulations, where large arrays of complex numbers are processed frequently.
II-B NUMA-Aware Optimization
Modern multi-socket servers are NUMA: each CPU has faster access to memory on its local socket than on remote sockets. If left to the default OS policies, the memory for a process may be spread across NUMA nodes in a suboptimal way, and threads may migrate between sockets, incurring costly remote memory accesses. Often this is controlled in the HPC setting by pinning MPI processes to NUMA regions, but this approach is sub-optimal in QuEST, due to the additional memory buffers associated with MPI communication increasing the memory overhead. Thereby, in this section, we focus on two strategies to optimize memory placement and thread affinity on a dual-NUMA system.
II-B1 Memory Allocation
By default, QuEST allocates the state-vector with the standard libc malloc, which on Linux follows a first-touch policy for NUMA (pages are allocated on the node where they are first accessed). This means that the OS is free to distribute pages across NUMA nodes arbitrarily, which may not be optimal on systems with multiple memory nodes, leading to performance penalties from remote memory accesses. To address this limitation, we introduce explicit control over memory placement via two NUMA-optimized strategies: (i) Local Allocation (V1), where the code first queries the free memory on each NUMA node (via numa_node_size64), then checks if the required memory can fit into a single node. If so, it performs an mmap allocation and sets a bind policy with mbind to restrict that allocation strictly to the chosen node. This ensures localized memory access – maximizing the benefit of a minimal memory latency. If the memory size exceeds what one node can provide, it gracefully falls back to partial allocation on the other node, while still enforcing page alignment. (ii) Split Allocation (V2), where we always split the allocated memory evenly across two nodes from the start. After computing halfSize (and ensuring that it remains at least one page in size), we allocate the entire region via mmap, then explicitly bind the first half of the region to Node 0 and the second half to Node 1. This approach ensures a balanced distribution of memory which yields higher total memory bandwidth. As in V1, the routine first rounds the requested buffer size up to an exact multiple of the system page size, allocates it with mmap, and verifies that the returned pointer is 64-byte aligned – guaranteeing efficient AVX-512 loads and stores. Both strategies confirm that the base address still starts on a page boundary, eliminating sub-page fragmentation that could otherwise hamper memory-access performance. Figure 2 compares the available allocation strategies schematically. These methods integrate seamlessly within the QuEST environment, allowing the user to switch between Default, NUMA V1, or NUMA V2 at compile or runtime, depending on the specific memory requirements and system architecture.
II-B2 Thread Affinity
Locality‐aware task scheduling is essential in NUMA architectures because it ensures that threads run on the memory node where their data reside, thereby minimizing costly remote accesses. In our approach, we dynamically calculate task chunks and bind threads to specific NUMA nodes. Instead of relying on the OS scheduler, we bind half of the threads to the first socket and the other half to the second socket – aligned with layout used in allocation V2 – see Algorithm 2. Moreover, we divide the state-vector workload such that each thread mostly works on the portion of the state-vector residing in its local node’s memory. In our two-node system, threads process the lower half of the amplitude array (allocated on node 0) and threads process the upper half (allocated on node 1). This manual partitioning and pinning maximizes data locality for most workloads and prevents idle resources when tasks are not perfectly divisible. Together with optimizations such as loop unrolling and cache‐friendly data layouts, this strategy significantly lowers memory‐access latency and boosts overall performance. Figure 3 evaluates the performance impact of NUMA-aware optimizations on a sequence of 50 Hadamard gates, profiled with Intel PCM NUMA [19]. The Y-axes show the DRAM accesses as seen by different threads (on the X-axes), for different qubits. Configuration (a) employs the baseline QuEST implementation, leading to suboptimal/scattered accesses across nodes. Configuration (b) introduces NUMA-aware memory allocation while still using the baseline Hadamard gate, reducing remote memory accesses and improving memory locality by preferentially utilizing local memory. Configuration (c) further refines this approach utilising a locality‐sensitive scheduler. As a result, when the working set fits entirely within a single NUMA node – up to 34 qubits on our test system – all DRAM accesses remain local, effectively minimizing remote memory traffic. When total memory demand spans multiple nodes (35 qubits), the state-vector spans NUMA nodes, resulting in a mix of local and remote DRAM accesses, thereby still minimizing cross‐node traffic. Moreover, similar improvements can be achieved by applying environment settings (e.g., OMP_PROC_BIND, OMP_PLACES, GOMP_CPU_AFFINITY) that enforce core affinity, confirming the critical role of thread binding in reducing latency.
II-C Instruction-level Tuning
With the data layout and NUMA strategy in place, we further optimize the innermost loop of the simulation using advanced instruction-level techniques.
II-C1 Vectorization with AVX-512 Intrinsics
Vectorization allows a processor to more efficiently exploit data-parallelism, by performing identical instructions on different data in a single issue cycle. This often leads to performance gains when applications exhibit this pattern of having the same operation to many operands. In the case of state-vector simulation, the core operation is a matrix-vector multiplication, which exhibits enormous data parallelism, given the known regularity of the matrix sparsity pattern. However, enabling compiler optimizations at the compiler is insufficient to see any effect. This is due to how the code is written. In particular, the core of the gate loops in QuEST v3.7 rely heavily on bitwise operations, in particular masking and shifting, to identify which iterations to skip in controlled gates, and to calculate accessed indices from the iteration index. This pattern that does not translate well to vectorised instructions. Additionally, even when vectorising is working as expected and instructions are issued out faster and more efficiently, the memory bandwidth bottleneck still prevents translating these gains in faster gate execution times. In our case, we explicitly write our optimized kernels using intrinsic AVX-512 instructions, which permit simultaneous operations on eight double-precision floating-point values (4 complex amplitudes). Using aligned vector loads and stores (via _mm512_load_pd and _mm512_store_pd), the code minimizes memory transfer overhead. Critical arithmetic is further optimized using FMA instructions (_mm512_fmadd_pd and _mm512_fmsub_pd), which fuse multiplication with addition or subtraction operations common in quantum gate updates, thus further reducing the instruction count and improving numerical accuracy.
II-C2 Loop Unrolling and Pointer Restriction
Aggressive loop unrolling is employed – processing blocks of 16 iterations at a time – to diminish the overhead of loop control and enhance instruction-level parallelism to the CPU. The use of the restrict qualifier further assists the compiler by confirming non-aliasing pointers, allowing more aggressive optimizations during vectorization.
II-C3 Memory Optimization and NUMA Awareness
In addition to vector-level improvements, the implementation incorporates memory optimization strategies that are critical for high-performance execution. Utilising OpenMP, each thread operates on a distinct segment of the state vector, and a NUMA-aware approach is adopted by binding threads to specific NUMA nodes (using numa_run_on_node) to mitigate cross-node memory latencies and optimize memory locality.
II-C4 Memory Prefetching
To further reduce memory access latency, for each group of 4 complex numbers processed, the code explicitly employs prefetching via _mm_prefetch to load future cache lines in advance, reducing cache miss penalties when processing large blocks of data. Our kernel processes 8 doubles at a time, hence prefetching data 16 and 32 elements ahead ensures that the subsequent iterations’ data is already loaded into the L1 cache.
II-C5 Scalar Fallback for Complete Coverage
To guarantee complete processing of the state vector, any iterations that cannot be handled by the vectorized loop are managed via a scalar fallback routine. This ensures that all amplitudes are accurately transformed, even when they do not align with the vectorized block sizes. A similar prefetching (with a smaller offset) ensures even the remaining, non-vectorized elements benefit from lower cache miss penalties.
These optimizations are illustrated in pseudocode: Algorithm 1 shows the baseline Hadamard transformation incorporating the new AoS, while Algorithm 2 shows our fully optimised kernel. This implementation is engineered to maximize throughput on modern multi-core architectures through a series of low-level optimizations presented in Section II.
III Benchmark Results and Analysis
The experiments were conducted on a dual-socket Intel® Xeon® Gold 6230R system (x86_64) featuring 52 physical cores (26 per socket), with hyperthreading disabled to ensure consistent core performance. Each core includes 32KiB instruction and 32KiB data L1 cache, and 1 MiB L2 cache. Each socket includes 35.75MiB L3 cache, for a total of 71.5 MiB amongst the two sockets. The system is configured with two NUMA nodes, each offering approximately 385 GB of local memory. Notably, local memory latency is about 100 ns, while remote memory latency reaches 150 ns, as measured using Intel’s MLC, highlighting the importance of NUMA-aware optimizations. To ensure stable and reproducible performance results, we disabled Intel Turbo Boost and turned off CPU frequency scaling, preventing transient clock speed variations from influencing the benchmarks. All experiments were performed on a Linux environment, and the code was compiled in Release mode with GCC 9.3.0 (C99, C++98) via CMake 3.7, employing aggressive optimization flags (e.g., -O3) to fully leverage the hardware’s performance capabilities. Moreover, for consistent benchmarking, we utilized Google Benchmark integrated within the Yao framework [20], where
to enhance statistical reliability each benchmark operation (i.e., an individual gate or a composite circuit) was repeated ten times and the minimum simulation time is reported.
III-A Single- and Two-Qubit Operations
Figure 6(a) reports the absolute execution time per gate for a representative single-qubit gate (Hadamard). The baseline QuEST results are compared against multiple optimization strategies, and the effectiveness of these optimizations is more evident in Figure 6(b), which shows the corresponding speedups relative to the baseline. The overall performance gains observed across our optimizations indicate the clear benefits of progressively more aggressive low-level tuning. Reordering the data into an AoS alone yields a modest speedup of 1.2× over the baseline. Adding AVX2 vectorization dramatically increases the speedup – 5× faster than baseline. When we switch from AVX2 to AVX-512 (with the AoS layout), the speedup improves further to 5.5×. The best performance is achieved by the most aggressive optimization (AoS + AVX-512 + FMA + loop unrolling + prefetching), see Section II-C and Algorithm 2, reaching a speedup in the vicinity of 6× compared to the baseline. These gains reflect how each low-level improvement builds upon the previous ones and can be directly attributed to the underlying memory-access patterns of quantum operations. The restructured AoS format enhances spatial locality, which in turn enables efficient SIMD loading of data.
For single‐qubit operations, the in‐place matrix‐vector multiply (depicted in Figure 4) follows a relatively regular and predictable access pattern that is well‐matched to cache‐prefetching and advanced SIMD optimizations (AVX2 and especially AVX-512). There is minimal branching in these kernels; every qubit undergoes a uniform sequence of single‐qubit transformations in parallel, leading to dense, vectorizable memory loads and stores. As every amplitude is processed and there is minimal branching, the kernel processing each target qubit traverses (and implicitly permutes) the entire state vector in a systematic fashion that maximizes cache reuse and vectorization across CPU cores – resulting in significant performance gains. Overall, our findings underscore that while both memory reorganization and vectorization contribute to improved performance, the combination of the new AoS with NUMA-aware tuning and instruction-level tricks is most effective in leveraging the regularity of the memory access pattern to yield the largest speedup when scaling simulations to higher qubit numbers.
Building on the optimizations introduced in Figure 6(b), Figure 7 extends this analysis to compare the three gate types (Hadamard, Pauli-, and CNOT) using the fully optimized kernel under the two NUMA memory allocation strategies (see Section II-B1). V1 (node‐0-first) remains faster than V2 (split allocation) up to the penultimate qubit size. In this region, the memory footprint fits comfortably within a single NUMA node, so the overheads of remote accesses are minimal, and local allocation yields lower latency. However, once we reach the largest qubit size (where one node’s memory no longer suffices), V1 is forced to allocate data on the remote node as well, causing its performance to degrade. At this point, V2 (which evenly distributes memory across both nodes) achieves comparable or better results by balancing the load between nodes. Another key observation is that speed-up improvements differ notably among the three gate types. For Hadamard and Pauli- operations, we achieve higher speedups (up to 5.5-6.5×) over the baseline compared to CNOT (which peaks around 4.5×). This behavior is closely tied to the memory‐access patterns depicted in Figures 5. In contrast to the single-qubit gates, which involve uniform, independent rotations on all amplitudes (ideal for vectorized execution), CNOT involves controlled operations, where certain branches of the matrix‐vector multiply are skipped depending on the control qubit’s state – only the amplitudes corresponding to a control bit of 1 are updated. This conditional behavior introduces branches in the matrix‐vector multiply, as certain computational paths are effectively skipped. This branching logic and the more irregular, conditional nature of the data manipulation reduces opportunities for continuous SIMD operations and hampers consistent cache reuse. Consequently, although CNOT still benefits from multi‐threading and low-level tuning strategies, it does not achieve speedups as high as the simpler single‐qubit operations.
III-B Quantum Fourier Transform
To evaluate our approach on a more complex multi-qubit algorithm, we examined the QFT circuit. Figure 8(a) compares the execution times of the baseline QFT across ten qubit sizes using four different implementations: a divide-and-conquer recursive variant inspired by libraries like Yao/Qiskit, an iterative method, QuEST Built-in, and an optimized cache-blocked version with two blocking factors. The recursive implementation mirrors the mathematical definition of the QFT by recursively applying a Hadamard gate followed by controlled phase shifts. The iterative method, on the other hand, rearranges the qubit operations into nested loops instead of relying on recursion. The iterative layout can be more cache-friendly in certain scenarios due to its straightforward memory access pattern, representing differences of less than 0.5–1.5% relative to the recursive approach across all qubit sizes. Although the built-in method is robust, it yields consistently higher execution times. At 23 qubits, the QuEST-Builtin runs 1.86× slower than the recursive/iterative variants, and at 32 qubits it’s roughly 1.70× slower. This indicates that the hand-optimized approaches can provide a significant performance gain – roughly a 70–80% reduction in runtime compared to the default library implementation. The QFT can be cache-blocked without extra gates by shifting its ending SWAP gates to the left, ensuring that all Hadamard gates operate on local qubits. Only the distributed SWAP gates (which are flipped vertically) incur communication overhead [15]. In an effort to improve data locality, we implemented a cache-blocking technique with a blocking factor of 16, which is 3% slower than the iterative approach at 23 qubits. At the upper end (32 qubits), the blocked variant is about 4.8% slower than the iterative version. These results suggest that for the chosen block size, the overhead associated with managing cache blocks slightly outweighs the potential gains in memory locality, particularly as the data size grows. When we increased the block size to 64, performance remained virtually identical to the iterative baseline – evidence that the iterative method already exploits strong memory locality. The larger block size simply further amortizes data‐movement costs and maximizes cache utilization without adding significant overhead.
To evaluate the effectiveness of our optimizations, we compare two implementations of QFT – recursive and QuEST Built-in – as summarized in Table II. Overall, the recursive variant consistently achieves higher speedups (1.8x) than the Built-in (1.7x), indicating that its code structure may benefit more from compiler‐driven optimizations like inlining or function‐call elimination, giving it a minor advantage. Nevertheless, both implementations show robust improvements over their respective baselines due to low‐level enhancements. Hence, while both implementations benefit from our optimizations, the recursive QFT’s design appears to leverage these techniques slightly more effectively. Nonetheless, both QFT implementations are less amenable to extreme speedups compared to simple gate operations. The QFT’s series of dependent rotations and swaps impose a more sequential data access pattern, which limits the gains achievable from SIMD and caching techniques.
| Qubit | Recursive QFT | Built-in QFT |
| 23 | 1.79 | 1.67 |
| 24 | 1.75 | 1.66 |
| 25 | 1.76 | 1.65 |
| 26 | 1.77 | 1.67 |
| 27 | 1.78 | 1.67 |
| 28 | 1.76 | 1.66 |
| 29 | 1.77 | 1.68 |
| 30 | 1.75 | 1.66 |
| 31 | 1.76 | 1.65 |
| 32 | 1.77 | 1.66 |
III-C Random Quantum Circuits
As a final benchmark, we implemented an RQC generator to stress-test the simulator. This setup allows us to quantify the impact of our low-level optimizations on both the dynamic generation of random single‑qubit gate sequences and the highly entangled execution of CNOTs connecting qubit →. Here, each circuit comprises a fixed number of layers (), and we sweep the register size from 23 to 32 qubits. We measured the performance of executing these circuits end-to-end, including state initialization, random gate generation, and state destruction, using the Google Benchmark framework for rigorous timing. Figure 9 presents the results for the baseline and optimized RQC executions. Despite the randomness of the circuits, their structure is fairly regular across qubits and layers, and our optimized simulator achieves a consistent speedup of 4× over the baseline. The RQC’s uniform layering and parallel gate structure are extremely well-suited to our low-level optimizations, which greatly reduce cache misses and overlap memory accesses with computation, thereby yielding a substantial boost in overall simulation speed. By contrast, as discussed above, the QFT’s inherently sequential structure and greater arithmetic complexity, which is primarily due to the interdependent controlled phase rotations introduces bottlenecks that constrain the effectiveness of SIMD strategies. Loop unrolling and prefetching do not capitalize on these operations as efficiently, and the sequential dependencies introduce communication overhead that further hinders performance improvements-limiting its speedup to 1.8×. This comparison highlights that the effectiveness of our optimizations can depend on the circuit’s structure – highly parallel circuits like RQCs reap the most benefit, whereas circuits with serial dependencies see more moderate gains.
IV Discussion
Scaling. Our experiments only include configurations with 2 NUMA nodes. While this is not necessarily the most NUMA regions that can be observed in modern machines, we expect our evaluation to cover an overwhelming amount of them nonetheless, for two main reasons. Firstly, when more compute power is needed, modern systems often reach for CPUs with more cores per socket, or accelerators such as GPUs or TPUs, rather than large socket counts per node. Machines that need a large number of CPUs or GPUs usually fall directly under the distributed computing paradigm, where shared memory is no longer applicable [21]. Secondly, in the presence of several NUMA regions on 2 sockets, the penalties paid for going across sockets are much larger than when going across memory controllers on the same socket. We observe for example on ARCHER2 [22], a machine running dual-socket AMD® EPYC 7742 processors nodes with 8 NUMA regions (4 per processor), that going across sockets in the same node costs a 30% decrease in memory bandwidth – similar to the performance of the system we used for benchmarking – while going across NUMA regions on the same socket costs 2% to 5%. This loss is so much smaller than going across sockets, that functionally this 8-region machine behaves like a 2-region configuration in practice.
Portability. Most elements of our optimisation stack are machine-agnostic. NUMA-aware page allocation – implemented with mmap + mbind – works on any Linux host that ships libnuma; when the library is missing or the system is single-socket, an #ifdef HAVE_LIBNUMA guard transparently falls back to plain mmap. Vectorisation is selected at compile-time: the fastest path uses AVX-512 intrinsics (__m512d, _mm512_*) available on recent Intel processors (Skylake-X, Cascade Lake, Ice Lake, Sapphire Rapids). For wider coverage we supply: (i) an AVX2 variant built with __m256d, (ii) a compiler-guided fallback that relies on #pragma omp simd when no explicit SIMD ISA is present, and (iii) automatic ISA dispatch via GCC/Clang function multiversioning (__attribute__((target("avx512f")))) or build-time flags such as -mavx512f, -mavx2, and -march=native, which lets the compiler pick the widest instruction set supported by the build host automatically. Thread scheduling stays portable because the loop is statically divided across cores, but the simple “first-half to node 0” binding must be replaced by a topology query on machines whose cores are not split 50/50. Finally, the two-cache-line prefetch distance is exposed as a tunable constant (PREFETCH_AHEAD 32) so it can be retuned for processors with different LLC sizes or memory latencies. In short, every layer that drives the bulk of the speed-up either ports unchanged or degrades gracefully, enabling consistent acceleration on both modern Intel and AMD servers.
V Conclusion
Our work bridges the gap between high-level algorithmic efficiency and low-level, architecture-specific performance enhancements. We have significantly enhanced the robust foundation of the QuEST simulator (v3.7) by developing a NUMA-aware, vectorized execution model optimized explicitly for modern multi-core processors. By restructuring the data layout to an AoS and applying targeted optimizations – including NUMA-aware memory allocation, explicit thread pinning, AVX-512 vectorization with intrinsic instructions, aggressive loop unrolling, and explicit memory prefetching – we achieved substantial speedups: 5.5–6.5× for single-qubit gates, 4.5× for two-qubit gates, 4× for RQCs, and approximately 1.8× for QFT circuits over QuEST. Our optimizations compare favorably against related works such as HpQC and PAS. HpQC similarly utilizes AoS memory layout and AVX2 intrinsics, reporting QFT speedups of 2.20× and RQC speedups around 1.74×. PAS incorporates distinct optimizations such as hybrid vectorization (combining intrinsic and compiler auto-vectorization), fast bitwise arithmetic, quantum tracking to avoid redundant computations, and memory access filtering, achieving higher performance gains – up to 8.69× on QFT and 2.62× on RQC. PAS also explicitly emphasizes large 4K-byte aligned memory allocations compared to our 64-byte alignment strategy, reflecting a more aggressive optimization for large memory pages. Despite these notable techniques in PAS and HpQC, our simulator uniquely integrates NUMA-aware memory placement and explicit core affinity, significantly reducing inter-node latency and better exploiting memory bandwidth on NUMA systems – features notably absent from PAS and HpQC. Moreover, unlike PAS and HpQC – which currently do not provide open access to their implementations – our enhancements are fully open-sourced, promoting transparency, reproducibility, and community-driven improvements. Our extension is hosted in our group’s repository [23], available upon reasonable request, and will soon be integrated into mainstream QuEST following completion of integration testing. This will enable a comprehensive benchmarking of our optimizations directly against the latest QuEST v4 release [24], which introduced new algorithmic methods [2] and architectural enhancements, but without incorporating our detailed NUMA-aware and SIMD-based tuning proposed here. Such benchmarking will quantitatively assess the specific performance impacts attributable to our low-level optimizations. Ultimately, our framework provides a modular, extensible foundation for accelerating classical quantum simulations.
Future Work. Going forward, we aim to maximize single-node performance and scalability through automation of vectorized kernel generation and compile-time-driven prefetching strategies, while extending support to larger shared-memory systems and heterogeneous accelerator-based platforms. In the near term, we plan to adapt NUMA-aware optimization concepts to GPUs, leveraging their shared emphasis on data locality. Although NUMA originated in CPU architectures, its core principles – memory locality and task affinity – are increasingly applicable to GPUs and can inform performance-aware execution across heterogeneous platforms [25]. Next, we will explore how emerging CXL memory expanders – i.e., CPU-less NUMA nodes impact the performance and applicability of the techniques introduced in this work. While CXL clearly extends the memory capacity of a single node, its effect on simulation performance and the strategies required to optimize it remain open questions. Finally, we will explore low-level and locality-sensitive challenges in distributed multi-node, multi-GPU environments that leverage MPI and RDMA for communication.
Acknowledgment
This work was supported by the EPSRC under grant EP/W032635/1 as part of the RoaRQ (Robust and Reliable Quantum) project. The authors thank the RoaRQ consortium for providing financial support and a collaborative environment that enabled the development of the methods and results presented in this paper.